I thought I had the answer but I'm stuck. Can you solve this?
A rectangular lot whose perimeter is 320 ft is fenced along three sides. An expensive fencing along the lot's length costs $20 per foot and an inexpensive fencing along the two side widths costs only $8 per foot. The total cost of the fencing along the three sides comes to $2140. What are the lot's dimensions?

Answer by scott8148(6628)

2L + 2W = 320 ___ L + W = 160 ___ 4L + 4W = 640

20L + 8W + 8W = 2140 ___ 5L + 4W = 535

subtracting the equations to eliminate W, leaves a negative value for L

it appears that the total fencing cost is awry

assuming a square lot (max area with min perimeter) ___ 80(20) + 80(8) + 80(8) = 2880